OpenStudy (christos):

if tan^2θ=8 find the exact value of sec^2θ

OpenStudy (anonymous):

http://bit.ly/QU4mW2

OpenStudy (anonymous):

45

OpenStudy (christos):

but how

hartnn (hartnn):

\(\sec^2x = 1+\tan^2 x\)

OpenStudy (anonymous):

Recall the Pythagorean identity:

OpenStudy (christos):

csc^2x = 1 + cot^2x too @hartnn ?

hartnn (hartnn):

yes.

OpenStudy (christos):

thanks!!!

OpenStudy (anonymous):

 By the Pythagorean theorem, we know \(a^2+b^2=c^2\), or, in this case, \(\cos^2x+\sin^2x=1\). Divide through by \(\cos^2x\) to yield \(1+\frac{\sin^2x}{\cos^2x}=\frac1{\cos^2x}\) -- or, equivalently, \(1+\tan^2x=\sec^2x\)